OpenStudy (anonymous):

I still don't get it! What's 9+10?

OpenStudy (anonymous):

9+10=21

OpenStudy (anonymous):

Its 19

OpenStudy (anonymous):

910

OpenStudy (lcnettech):

9 + 10 = 19

OpenStudy (anonymous):

Use your fingers and toes

OpenStudy (anonymous):

its in a movie

OpenStudy (lcnettech):

just because it's in a movie doesnt mean that 9+10=21 it equals 19 .

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

BUT according to The Vine Academy it 21 But anyone with half a brain knows it 19 :D

OpenStudy (k_lynn):

Ok guys. Please don't respond anymore. Most of the mods consider the "9+10" post spam. Responding makes more spam. That's why I'm asking not to reply anymore. Report and move on.

HanAkoSolo (jamierox4ev3r):

this is such an old joke. Please stop. We don't want to spam the math question with non-serious questions :)

TheSmartOne (thesmartone):

Welcome to OpenStudy! @Gunboss Do you know that OpenStudy has a 13+ age policy? Wait till you are 13 or older to use this site. :)